OpenStudy (anonymous):

Find a point on the curve defined by function y = x^2 that has the minimum distance to the line defined by y = 2x - 4. (c) Solve this problem using Lagrange multiplier approach.

OpenStudy (anonymous):

When I tried solving this problem, I ended up with (1, -2), but I think the correct answer is (1,1).

OpenStudy (freckles):

You used the distance formula?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

When I used the Lagrange method, I got something different. Not sure how I made an error.

OpenStudy (freckles):

Did you have something like this: \[f(x_1,x_2)=(x_1-x_2)^2+(x^2_1-2x_2+4)^2 \] then we have \[f_{x_1}=\lambda g_{x_1 } \\ f_{x_2}=\lambda g_{x_2}\] but I'm having trouble on what the constraint is ... Can I see what you did?

OpenStudy (anonymous):

x^2 -y - lambda(2x-y-4) = F(x,y, lambda) x^2 - y -2lambda x + lamda y + 4 lamda partial x: 2x - 2lambda = 0 partial y: -1 + lamda = 0 partial lamda: -2x + y + 4 = 0. And then I just solved.

OpenStudy (freckles):

oh I think you just got it backwards

OpenStudy (freckles):

\[2x-y-4-\lambda(x^2-y)=F \\ F_x=2-2x \lambda=0 \\ F_y=-1+\lambda=0 \\ F_\lambda=-(x^2-y)=0\]

OpenStudy (freckles):

that works